State if each set of three numbers can be the lengths of the sides of a
triangle.
1. 5, 13, 17
2. 13, 4, 19
3. 13, 5, 15
4. 13, 5, 13
5. 8, 3, 7

Answer by greenestamps(13367)



Three lengths can make a triangle if the sum of the two smallest numbers is greater than the largest.

Example 1:
8,10,12 YES: 8+10=18 is greater than 12
Example 2:
4,7,12 NO: 4+7=11 is less than 12
Example 3:
5,6,11 NO: 5+6=11 is not GREATER THAN 11

You can figure out the answers for each set of lengths in this problem.